OpenStudy (anonymous):

Use the method of substitution to solve the system of linear eqautions. 2x-y=6, 9x-4y=28. The solution is____

OpenStudy (darkprince14):

express 2x-y=6 in terms of either x or y... 2x - y = 6 (transpose y) 2x = 6 + y (divide both sides by 2 to know what is x) x = 3 + y/2 or 2x - y = 6 (transpose 2x) -y = 6 - 2x (multiply -1 both sides) y = 2x - 6 Then pud it into the other equation...

OpenStudy (darkprince14):

*plug